# Wave–Particle Complementarity as the Optimal Limit of Unambiguous Quantum-State Discrimination

### v2

@@ C.1 Orthogonal complements of the idler states
C.2 POVM elements and zero-error conditions
@@
C.3 Success and inconclusive probabilities
@@
C.4 Positivity of $\Pi_{0}$ and optimal coefficients
@@
C.5 Relation to visibility and complementarity
From Appendix A, the ZWM signal visibility is
@@ C.5 Relation to visibility and complementarity
Appendix D Schrödinger-picture derivation of the two-crystal ZWM state
@@
Appendix E General two-path interferometer with which-path markers
@@
E.1 A. Joint state and reduced density matrix
@@
E.4 D. IDP optimal failure probability
The IDP strategy for unambiguous discrimination between the two pure marker states $|\mu_{1}\rangle,|\mu_{2}\rangle$ with priors $\eta_{1},\eta_{2}$ and overlap $\gamma$ achieves an optimal inconclusive probability [Ivanovic, Dieks, Peres, Herzog2005MixedUSD]
$P_{I}^{\mathrm{opt}}=2\sqrt{\eta_{1}\eta_{2}}\,|\gamma|,$ (114)
and an optimal success probability
$P_{S}^{\mathrm{opt}}=1-P_{I}^{\mathrm{opt}}=1-2\sqrt{\eta_{1}\eta_{2}}\,|\gamma|.$ (115)
Comparing with Eq. (108), we see that
$P_{I}^{\mathrm{opt}}=V.$ (116)
Hence the distinguishability can be expressed as
$D^{2}=1-V^{2}=1-\big(P_{I}^{\mathrm{opt}}\big)^{2},$ (117)
which yields the operational relation
$D^{2}+\big(P_{I}^{\mathrm{opt}}\big)^{2}=1.$ (118)
E.5 E. Remarks on unequal priors and mixed markers
@@
Appendix F Experimental protocol
@@
Appendix G Mixed idler marker states with thermal noise
In the main text we worked in the ideal low-gain regime with pure idler marker states, for which the overlap $|\langle\phi_{1}|\phi_{2}\rangle|$ fixes both the single-photon visibility and the optimal IDP failure probability, $V=|\langle\phi_{1}|\phi_{2}\rangle|=P_{I}^{\mathrm{opt}}$. In this Appendix we outline how the picture is modified when thermal background photons are injected into the idler path, so that the which-crystal markers become mixed states. We work in the Schrödinger picture at the level of density operators, and make use of Uhlmann fidelity rather than repeating the full Heisenberg calculation [boyd, giri1].
G.1 A. Thermal environment and conditional idler states
@@ Appendix G Mixed idler marker states with thermal noise
G.2 B. Purifications, visibility, and Uhlmann fidelity
@@
G.3 C. Optimal IDP failure probability for mixed markers

### v3

@@ E.1 Orthogonal complements of the idler states
E.2 POVM elements and zero-error conditions
@@
E.3 Success and inconclusive probabilities
@@
E.4 Positivity of $\Pi_{0}$ and optimal coefficients
@@
E.5 Relation to visibility and complementarity
From Appendix C, the ZWM signal visibility is
@@ E.5 Relation to visibility and complementarity
Appendix F Schrödinger-picture derivation of the two-crystal ZWM state
@@
Appendix G General two-path interferometer with which-path markers
@@
G.1 A. Joint state and reduced density matrix
@@
G.4 D. IDP optimal failure probability
[…]
$P_{\rm inc}^{\mathrm{opt}}=2\sqrt{\eta_{1}\eta_{2}}\,|\gamma|,$ (178)
and therefore, comparing with Eq. (172),
$P_{\rm inc}^{\mathrm{opt}}=V\qquad\text{(symmetric-IDP branch)}.$ (179)
Remark on unequal priors.
For unequal a priori probabilities, the optimal unambiguous state discrimination (USD) strategy is in general piecewise. The expression (178) holds in the symmetric-IDP regime, which includes the equal-prior case $\eta_{1}=\eta_{2}=\tfrac{1}{2}$ relevant to equal pumping in the ZWM interferometer. Outside this regime, the optimal solution changes branch, and one of the conclusive outcomes may be suppressed. Since the present work focuses on the equal-pump induced-coherence scenario, we restrict attention to the symmetric-IDP branch.
G.5 E. Remarks on unequal priors and mixed markers
@@
Appendix H Experimental protocol
Previous unambiguous path discrimination in Mach–Zehnder interferometers [Len2018IDP, Clarke2001IDP], where USD identifies orthogonal signal paths via an auxiliary degree of freedom. The present work applies unambiguous state discrimination to a fundamentally different task. In the low-gain induced-coherence interferometer, the idler modes $I$ are deliberately aligned to be path-identical, rendering which-path information ill-defined for the idler photon. The only meaningful alternative encoded in the idler is therefore the source of the photon pair, namely which nonlinear crystal emitted it. Emission from crystal A and crystal B prepares two distinct but generally nonorthogonal conditional idler states. Unambiguous state discrimination on the idler thus constitutes a zero-error retrodictive measurement of the emitting crystal, rather than a which-path measurement. This distinction is essential: USD here addresses source identification in an induced-coherence geometry, not path discrimination in an interferometer.
@@ Appendix H Experimental protocol
Appendix I Mixed idler marker states with thermal noise
In the main text we worked in the ideal low-gain regime with pure idler marker states, for which the overlap $|\langle\phi_{1}|\phi_{2}\rangle|$ fixes both the single-photon visibility and the optimal IDP failure probability, $V=|\langle\phi_{1}|\phi_{2}\rangle|=P_{I}^{\mathrm{opt}}$. In this Appendix we outline how the picture is modified when thermal background photons are injected into the idler path $S$. The which-crystal markers become mixed states. We work in the Schrödinger picture at the level of density operators, and make use of Uhlmann fidelity rather than repeating the full Heisenberg calculation [boyd, giri1].
I.1 A. Thermal environment and conditional idler states
@@ Appendix I Mixed idler marker states with thermal noise
I.2 B. Purifications, visibility, and Uhlmann fidelity
@@
I.3 C. Optimal IDP failure probability for mixed markers